# Secure Smart Contract with Control Flow Integrity

### v1

@@ 5. Approach
In this section, we present the design of our control flow integrity technique. We first introduce the whitelisting policies. We then present a series of control flow simplification heuristics designed to increase the likelihood of the corresponding transaction meeting the whitelisting criteria. Lastly, we describe how these policies and heuristics are implemented within a smart contract monitor and how this monitor is applied to track transactions across different scenarios.
@@ 5.1. Control Flow Whitelisting Policies
We propose four control flow whitelisting policies to identify benign invocations.
Policy 1: Simple Independent Invocations. An invocation in a transaction is considered benign if it is simple and independent of any prior invocations executed within the transaction. The rationale for this policy is that a simple, independent invocation mirrors the behavior of a function being invoked by an EOA in a single, standalone call. This represents the fundamental usage of a function  [^1] If invoking a single function without re-entrancy within $P$ is flagged as malicious, it points to a single function access control issue, which falls outside the scope of this paper.. As such, this type of invocation is considered benign.
Policy 2: Read-Only Invocations. An invocation is considered benign if it is read-only and does not modify any blockchain state. Specifically, an invocation is read-only if it performs no write operations to the storage (i.e., no SSTORE operations), does not call other state-changing functions, and does not transfer Ether  [^2] Other operations, such as SELFDESTRUCT or CREATE, could also alter the blockchain state but are rare in high-profile DeFi protocols. If such operations are present in a function, that function should not be classified as read-only.. This is because invocations which do not alter the blockchain state have no effect on the final state. Therefore, those invocations can be omitted from the transaction without influencing the overall outcome.
Policy 3: Runtime Read-Only (RR) Function Calls. A function call is considered runtime read-only if it does not have a storage write (SSTORE), and it performs no Ether transfers. Some functions may not be marked as read-only in their source code but behave as such on the fly. An invocation is runtime read-only if all function calls within it are runtime read-only. By tracking runtime behavior and identifying such function calls and invocations, we can prune these invocations from the control flow, as they do not alter the blockchain state.
Policy 4: Restore-on-Exit (RE) Storage Writes. This heuristic permits to safely ignore storage writes that temporarily alter values but restore the original state at the end of execution, i.e., the value returned in the first read operation (SLOAD) equals the one written by its last write operation, with no write preceding the first read. This increases the likelihood of classifying function calls as runtime read-only. A typical example is the re-entrancy guard pattern, where a function restores the original state of the guard before exiting to prevent re-entrancy attacks. In our system, while these re-entrancy guards remain active and function as intended, any storage writes to them are disregarded when determining whether a function call is runtime read-only or not.
@@ 5.2. Control Flow Simplification Heuristics
To reduce the complexity of control flows and minimize the risk of flagging benign transactions as malicious, we propose two heuristics that allow CrossGuard to safely ignore certain function calls.
Heuristic 1: ERC20 Function Calls. ERC20 is a widely used smart contract standard for implementing tokens in DeFi protocols. ERC20 contracts perform token management, and several functions within these contracts modify user properties without impacting the overall protocol state. For example, the functions transfer and transferFrom change only the balances of the sender and receiver, while approve, increaseAllowance, and decreaseAllowance simply modify the allowance granted by the sender to the spender. We consider these five ERC20 functions to be benign and safe, as they have been extensively tested and are widely used across numerous DeFi projects. Therefore, calls to those functions within invocations are safely ignored.
Heuristic 2: Read-After-Write (RAW) Dependency. An invocation $\iota_{2}$ is considered storage read-after-write (RAW)-dependent on an earlier invocation $\iota_{1}$ if: $\iota_{2}$ reads from a storage location that $\iota_{1}$ writes to (with an exception of storage writes classified in Heuristic 1). In the absence of such dependencies, simple invocations are treated as independent, and a control flow consisting only of such simple and independent invocations is whitelisted. Note that re-entrant invocations, no matter it is dependent or not, will never be whitelisted by this heuristic.
@@ 5.3. System Overview
In this section, we explain how CrossGuard is integrated into a DeFi protocol pre-deployment by instrumenting the original code. The system consists of two main components: instrumentation within the protected contracts and a guard contract. Each function in the protected contracts is instrumented and assigned a unique positive integer as its function identifier.
@@
19:         $callTrace.push(-funcID)$
20:     if $sum=0$ then
21:         for each $id$ in $callTrace$ do
22:              $CFHash\leftarrow\text{keccak256}(id,CFHash)$
23:         $callTrace.clear()$
24:     if isRAW then
25:         $isCFRAW\leftarrow true$
26:     if $\neg$ _$\_allowedPatterns[CFHash]\land(isCFReEntrancy\lor isCFRAW)$ then
27:         revert “Unsafe pattern detected”
@@
When an instrumented function is invoked, it sends the guard contract its identifier to record its function entry and receives a positive integer value indicating the invocation count. The instrumented function then tracks storage accesses, records storage writes using the invocation count, and sends control flow data to the guard contract upon exit. The guard contract collects this control flow data from the protected contracts, simplifies it using the defined heuristics, evaluates whether the control flow is whitelisted, and reverts the transaction if it is not.  [^3] Note when protocol administrators execute a transaction, CrossGuard includes a straightforward mechanism (not detailed but trivial to implement) that allows administrators to deactivate the lock contract at the beginning of a transaction, perform actions without interference from the control flow integrity checks, and reactivate the lock contract at the end. Algorithm 1 outlines the implementation within the guard contract, while Algorithm 2 outlines a detailed description of both the execution and instrumentation applied to the protected functions. Furthermore, Table 1 specifies the instrumentation applied to the original source code.
@@
Algorithm in the guard contract. Algorithm 1 implements the control flow tracking in the guard contract. The function EnterFunc is invoked by protected contracts when one of their functions is called by an untrusted contract, taking a unique function identifier (funcID) as input for each function in the protected contracts. If the sum of function identifiers is zero, it signifies the start of a new invocation, and invCount is incremented (line 8). Otherwise, it indicates a re-entrancy condition, and isCFReEntrancy is set to true (line 10). The sum and callTrace are updated (lines 11-12) to record the function ID. The EnterFunc returns the invCount to the protected contracts for their future processing.
@@
The ExitFunc (lines 14-27) 444 Both EnterFunc and ExitFunc have access control to check if the caller is a protected contract, which is omitted for brevity in Algorithm 1. is called by the protected contracts at the exit of the same function that triggered EnterFunc. It takes three arguments: funcID, isRR (isRuntimeReadOnly), and isRAW (isRead-After-Write dependent on a previous invocation). If the invocation is runtime read-only, it removes the funcID added by EnterFunc from the callTrace (line 17). Otherwise, it pushes the negated funcID onto the stack (line 19). When the sum equals zero, signaling the end of an invocation, the CFHash is computed over the entire callTrace (lines 21-22) to summarize the control flow of the invocation as a hash. Additionally, if any invocation has a RAW dependency, the algorithm sets the isCFRAW flag to true (lines 24-25). Finally, if the computed hash does not match an allowed pattern, and a read-after-write condition or a re-entrancy condition is detected, the transaction is reverted to prevent unsafe behavior, enforcing the policy to block malicious control flows (lines 26-27). Without any pre-approved control flow patterns, _allowedPatterns only include simple invocations by default. But administrators can add more patterns to this mapping to whitelist more control flows.
@@
9:     for each $slot$ in $writeElements$ do
@@
13:     for each $slot$ in $readElements$ do
14:         for $i\leftarrow 1$ to invNum do
15:              if $storageWrites[slot,i]$ then
@@
18:     clear $tempReads$ and $tempWrites$
@@
After every SSTORE ($sstore(slot,value)$): 1:$writeElements.append(slot)$ 2:$tempWrites[slot]\leftarrow value$
@@
After every other state-changing opcode or external call to a state-changing function: 1:Set $isRR\leftarrow\textbf{false}$
After every call to other protected contracts ($funcCall()\to isSubRR,isSubRAW$): 1:$isRR\leftarrow isRR\land isSubRR$ 2:$isRAW\leftarrow isRAW\lor isSubRAW$
Algorithm for Protected Functions. Algorithm 2 implements the storage access tracking within instrumented protected functions. Table 1 provides a detailed breakdown of the instrumentation made to the original functions. Given an original function implementation, Func, it is replicated into two functions: FuncTrusted and FuncUntrusted. FuncTrusted can only be invoked by protected contracts 555 FuncTrusted has access control to check whether the msg.sender is in the whitelist of protected contracts, which is not detailed in Algorithm 2 for simplicity. , where the invNum is passed by its caller. In contrast, FuncUntrusted is designed to handle invocations from untrusted contracts or external wallets. It fetches the invNum from the guard contract by calling EnterFunc, tracks the storage accesses, and determines whether the function is runtime read-only or involves RAW (read-after-write) dependencies before sending the function ID to the guard contract via ExitFunc. Both functions call ExecuteInstrumentedCode, which tracks state changes and evaluates whether the invocation is runtime read-only and whether any read-after-write dependencies exist.
The algorithm initializes readElements and writeElements arrays to store accessed storage slots, alongside two boolean flags, isRR (runtime read-only) and isRAW (read-after-write dependencies), as described in lines 5-7 of the implementation. Then the algorithm proceeds to execute the instrumented code (line 8), with specifics provided in Table 1. During each SLOAD operation, instrumentation appends the accessed slot to readElements and records it in tempReads if it hasn’t previously been written to. Correspondingly, each SSTORE operation results in the slot being added to writeElements and its value stored in tempWrites. If any EVM opcode or function call modifies the blockchain state or if a subsequent protected contract call is not runtime read-only, isRR is set to false. Additionally, if any subsequent call to protected contracts involves a read-after-write dependency, isRAW is set to true.
@@
Finally, the algorithm checks and updates isRR and isRAW, as well as the storageWrites variable for future invocations (lines 9-19). For each written storage slot, the slot is recorded in storageWrites (line 10). If a slot written is not restored-on-exit, the function is marked as not runtime read-only(lines 11-12). For each slot read, the algorithm checks whether the slot was written to in a previous invocation, marking the invocation as RAW-dependent if so (lines 13-17). The temporary mappings are cleared (line 18) before returning isRR and isRAW to the guard contract (line 19).
5.4. A Running Example
Figure 2 illustrates how CrossGuard integrates with a series of protected contracts to detect and prevent malicious transactions initiated by attackers. In this scenario, the attacker uses a hack contract, referred to as Contract 0 (not shown in Figure 2 for simplicity), to launch an attack on the system. The attack begins when Contract 0 invokes FuncAUntrusted in Contract 1. This action triggers EnterFunc(funcID = 1) within the guard contract, which logs the function identifier and returns invCount = 1. Since Contract 1 is a trusted contract, it proceeds to call FuncBTrusted in Contract 2, a call permissible between trusted contracts without further interaction with the guard contract. Contract 2 uses the invCount passed from Contract 1 to monitor its read-only and read-after-write dependencies. After determining that its operations are runtime read-only and lack read-after-write dependencies, Contract 1 invokes ExitFunc(1, isRR = true, isRAW = false) in the guard contract. The guard contract, recognizing the operation as runtime read-only, removes the function call from the callTrace, effectively clearing it.
Figure 2. A Running Example of CrossGuard
Subsequently, the attacker directs Contract 0 to invoke FuncCUntrusted in Contract 3, leading to a call to EnterFunc(funcID = 2) by the guard contract, which assigns invCount = 2. Contract 3, under manipulation, then triggers an interaction with Contract 4—an untrusted contract—which in turn reenters Contract 3, causing it to execute FuncDUntrusted. This interaction prompts another call to EnterFunc(funcID = 3) in the guard contract. Here, the guard contract detects potential re-entrancy (setting isCFReEntrancy to true as outlined in line 10 of Algorithm 1) and records the function identifier, returning invCount = 3. Upon completion of its operations, Contract 3 calls ExitFunc(3, isRR = false, isRAW = true). Equipped with detailed logs of the control flow and detected violations (re-entrancy and read-after-write dependencies), the guard contract opts to revert the entire transaction. This decision is taken to block the execution of the malicious operation, effectively safeguarding the blockchain network from this and similar exploits. This robust mechanism provided by CrossGuard enables real-time detection and validation of advanced threats such as re-entrancy, ensuring control flow integrity across multiple protected contracts.
5.5. Optimizations
We have implemented two optimizations to reduce gas costs.
@@ 5.5. Optimizations
Optimization 1: Bypassing Validation for Simple Invocations from EOAs. When a protected contract’s function does not contain a re-entrancy bug and is directly invoked by an EOA, the transaction will consistently follow a single, straightforward invocation path, which conforms to the criteria set by Policy 1 (Section 5.1). We utilized Slither (Feist et al., 2019) to identify such functions. To account for the occasional false positives reported by Slither, we also conducted a manual verification of each function. A preliminary check is implemented at the beginning of these functions. If such functions are invoked by an EOA, the control flow validation can be safely bypassed, significantly reducing gas overhead.

### v2

@@ 5. Approach
In this section, we present the design of CrossGuard. We begin by introducing four whitelisting policies, followed by two heuristics aimed at reducing false positives. Finally, we describe how these policies and heuristics are implemented within a smart contract, and how they enable real-time control flow tracking and simplification.
@@ 5.1. Control Flow Whitelisting Policies
We propose four control flow whitelisting policies.
Policy 1: Simple Independent Invocations. An invocation in a transaction is considered benign if it is simple and independent of any prior invocations executed within the transaction. The rationale for this policy is that a simple, independent invocation mirrors the behavior of a function being invoked by an EOA in a single, standalone call. This represents the fundamental usage of a function  [^1] If a simple invocation is flagged as malicious, it points to a single function access control issue, which falls outside the scope of this paper.. As such, this type of invocation is considered benign.
Policy 2: Read-Only Invocations. An invocation is considered benign if it is read-only and does not modify any blockchain state. Specifically, an invocation is read-only if it performs no write operations to the storage (i.e., no SSTORE operations), does not call other state-changing functions, and does not transfer Ether  [^2] Other operations, such as SELFDESTRUCT or CREATE, could also alter the blockchain state but are rare in high-profile DeFi protocols. If such operations are present in a function, that function should not be classified as read-only.. This is because invocations not altering blockchain state have no effect on the final state. Therefore, those invocations can be omitted from the transaction without influencing the overall outcome.
Policy 3: Runtime Read-Only (RR) Function Calls. A function call is considered runtime read-only if it does not have a storage write (SSTORE), and it performs no Ether transfers. Some functions may not be marked as read-only in their source code but behave as such on the fly. An invocation is runtime read-only if all function calls within it are runtime read-only. By tracking runtime behavior and identifying such function calls and invocations, we can prune these invocations from the control flow, thereby simplifying it.
Policy 4: Restore-on-Exit (RE) Storage Writes. This heuristic permits to safely ignore storage writes that temporarily alter values but restore the original state at the end of execution, i.e., the value returned in the first read operation (SLOAD) equals the one written by its last write operation, with no write preceding the first read. This increases the likelihood of classifying function calls as runtime read-only. A typical example is re-entrancy guard, where a function restores the original state of the guard before exiting to prevent re-entrancy attacks. In our system, while these re-entrancy guards remain active and function as intended, any storage writes to them are disregarded when determining whether a function call is runtime read-only.
@@ 5.2. Control Flow Simplification Heuristics
To reduce the complexity of control flows and minimize false positives, we propose two heuristics that allow CrossGuard to safely ignore certain function calls.
Heuristic 1: ERC20 Function Calls. ERC20 (28) is a widely used smart contract standard for implementing tokens in DeFi protocols. ERC20 contracts perform token management, and several functions within these contracts modify user properties without impacting the overall protocol state. For example, the functions transfer and transferFrom change only the balances of the sender and receiver, while approve, increaseAllowance, and decreaseAllowance simply modify the allowance granted by the sender to the spender. We consider these five ERC20 functions to be benign and safe, as they have been extensively tested and are widely used across numerous DeFi projects. Therefore, calls to those functions within invocations are safely ignored.
Heuristic 2: Read-After-Write (RAW) Dependency. An invocation $\iota_{2}$ is considered storage read-after-write (RAW)-dependent on an earlier invocation $\iota_{1}$ if: $\iota_{2}$ reads from a storage location that $\iota_{1}$ writes to (with an exception of storage writes classified in Policy 4). In the absence of such dependencies, invocations are treated as independent, and a control flow consisting only simple and independent invocations is whitelisted. Note that re-entrant invocations, no matter if it is dependent, will never be whitelisted by this heuristic.
5.3. Soundness and Limitation Analysis
As mentioned in Section 4, CrossGuard is a general-purpose defense designed to prevent all smart contract attacks, with the only exception of those exploitable via a single function call to a protected contract (e.g., integer overflow or access control problems).
Policies 1-4 are sound by design as they only whitelist control flows that preserve security properties: Policy 1 ensures whitelisted transactions are equivalent to direct EOA calls, Policies 2-3 are inherently safe since read-only operations cannot alter blockchain state, and Policy 4 maintains soundness by ensuring temporary state changes are reverted, still not altering blockchain states.
However, our heuristics operate on a best-effort basis with specific limitations. Heuristic 1 assumes standard ERC20 functions maintain expected behavior, which could be violated by non-standard implementations. Heuristic 2 tracks read-after-write dependencies only within protected contracts and may miss external state manipulations such as oracle manipulations. However, as shown in Section 6, this limitation impacted only 1 out of 37 benchmarks evaluated (i.e. Bedrock_DeFi).
@@ 5.4. System Overview
In this section, we explain how CrossGuard is integrated into a DeFi protocol pre-deployment by instrumenting the original code. The system consists of two main components: instrumentation within the protected contracts and a guard contract. Each function in the protected contracts is instrumented and assigned a unique positive integer funcID as its identifier.
@@
19:   callTrace.push(funcID)callTrace.push(-funcID)
20:  $CFHash\leftarrow 0$
21:  if $sum=0$ then
22:   for each idid in callTracecallTrace do
23:     $CFHash\leftarrow\text{keccak256}(id,CFHash)$
24:   $callTrace.clear()$
25:  if isRAW then
26:   $isCFRAW\leftarrow true$
27:  if $\neg$ $allowedPatterns[CFHash]\land(isCFReEntrancy\lor isCFRAW)$ then
28:   revert “Unsafe pattern detected”
@@
When an instrumented function is invoked, it sends the guard contract its funcID to record its function entry and receives a positive integer invCount indicating the invocation count. The instrumented function then tracks storage accesses, records storage writes using invCount, and sends tracked runtime properties to the guard contract upon exit. The guard contract collects control flows and tracked runtime properties from the protected contracts, simplifies control flows using the defined policies and heuristics, evaluates whether the control flow is whitelisted, and reverts the transaction if it is not.  [^3] Note when protocol administrators execute a transaction, CrossGuard includes a straightforward mechanism (not detailed but trivial to implement) that allows administrators to deactivate the guard contract at the beginning of a transaction, perform actions without interference from the control flow integrity checks, and reactivate the guard contract at the end. Alg. 1 outlines the implementation within the guard contract, while Alg. 2 outlines the execution of the instrumented protected functions. Furthermore, Table 1 details the instrumentation applied to the original source code.
@@
Algorithm in the guard contract. Alg. 1 implements the control flow tracking in the guard contract. The function EnterFunc is invoked by protected contracts when one of their functions is called by an untrusted contract, taking a unique function identifier (funcID) as input for each function in the protected contracts. If the sum of function identifiers is zero, it signifies the start of a new invocation, and invCount is incremented (line 8). Otherwise, it indicates a re-entrancy condition, and isCFReEntrancy is set to true (line 10). The sum and callTrace are updated (lines 11-12) to record funcID. The EnterFunc returns invCount to the protected contracts, enabling them to internally track runtime properties.
@@
The ExitFunc (lines 14-28)  [^4] Both EnterFunc and ExitFunc have access control to check if the caller is a protected contract, which is omitted for brevity in Alg. 1. is called by the protected contracts at the exit of the same function that triggered EnterFunc. It takes three arguments: funcID, isRR (isRuntimeReadOnly), and isRAW (isRead-After-Write dependent on a previous invocation). If the invocation is runtime read-only, it removes the funcID added by EnterFunc from the callTrace (line 17). Otherwise, it pushes the negated funcID (which represents the end of the function call) onto the stack (line 19). When the sum equals zero, signaling the end of an invocation, the CFHash is computed over the entire callTrace (lines 22-23) to summarize the control flow of the invocation as a hash. Additionally, if any invocation has a RAW dependency, Alg. 1 sets the isCFRAW flag to true (lines 25-26). Finally, if the computed hash does not match an allowed pattern, and a read-after-write condition or a re-entrancy condition is detected, the transaction is reverted to prevent unsafe behavior, enforcing the policy to block malicious control flows (lines 27-28). Without any pre-approved control flow patterns, _allowedPatterns only include simple invocations by default. But administrators can add more patterns to this mapping to whitelist more control flows.
@@
9:  for each slotslot in writeElementswriteElements do
@@
13:  for each slotslot in readElementsreadElements do
14:   for $i\leftarrow 1$ to invNum1invNum-1 do
15:     if $storageWrites[i][slot]$ then
@@
18:  clear tempReadstempReads and tempWritestempWrites
@@
Algorithm for Protected Functions. Alg. 2 implements the storage access tracking within instrumented protected functions. Table 1 provides a detailed breakdown of the instrumentation made to the original functions. Given an original function implementation, Func, it is replicated into two functions: FuncTrusted and FuncUntrusted. FuncTrusted can only be invoked by protected contracts  [^5] FuncTrusted has access control to check whether the msg.sender is in the whitelist of protected contracts, which is not detailed in Alg. 2 for simplicity. , where the invNum is passed by its caller. In contrast, FuncUntrusted is designed to handle invocations from untrusted contracts or external wallets. It fetches the invNum from the guard contract by calling EnterFunc, tracks the storage accesses, and determines whether the function is runtime read-only or involves RAW (read-after-write) dependencies before sending the function identifier to the guard contract via ExitFunc. Both functions call
ExecuteInstrumentedCode, which tracks state changes and evaluates whether the invocation is runtime read-only and whether any read-after-write dependencies exist.
Alg. 2 initializes readElements and writeElements arrays to store accessed storage slots, alongside two boolean flags, isRR (runtime read-only) and isRAW (read-after-write dependencies), as described in lines 5-7 of the implementation. Then Alg. 2 proceeds to execute the instrumented code (line 8), with specifics provided in Table 1. After each SLOAD operation, instrumentation appends the accessed slot to readElements and records it in tempReads if it hasn’t previously been written to. Correspondingly, each SSTORE operation results in the slot being added to writeElements and its value stored in tempWrites. If any EVM opcode or function call modifies the blockchain state or if a subsequent protected contract call is not runtime read-only, isRR is set to false. Additionally, if any subsequent call to protected contracts involves a read-after-write dependency, isRAW is set to true.
@@
[…]
@@
After other state-changing opcodes or external non-read-only calls: 1:Set $isRR\leftarrow\textbf{false}$
[…]
@@
Finally, Alg. 2 checks and updates isRR and isRAW, as well as storageWrites for future invocations (lines 9-19). For each written storage slot, the slot is recorded in storageWrites (line 10). If a slot written is not restored-on-exit, the function is marked as not runtime read-only(lines 11-12). For each slot read, Alg. 2 checks whether the slot was written to in a previous invocation, marking the invocation as RAW-dependent if so (lines 13-17). The temporary mappings are cleared (line 18) before returning isRR and isRAW to the guard contract (line 19).
5.5. Gas Optimizations
We have implemented four optimizations to reduce gas costs.
@@ 5.5. Gas Optimizations
Optimization 1: Bypassing Validation for Simple Invocations from EOAs. When a protected function is not designed to invoke arbitrary untrusted contracts given by users( a prerequisite for re-entrancy attacks), and is directly invoked by an EOA, the transaction will consistently follow a single, straightforward invocation path, which conforms to the criteria set by Policy 1 (Section 5.1). We use Slither (Feist et al., 2019) to identify these functions and manually verify to mitigate occasional false positives. For the verified functions, we insert an EOA check at the beginning of execution. When the caller is an EOA, control-flow validation can be safely bypassed, substantially reducing gas overhead.
@@
Optimization 3: Merging Guard Contract. We merge the guard contract with the most frequently used protected contract to convert external EnterFunc and ExitFunc calls into internal calls, reducing gas overhead. Since it is hard to predict usage patterns before deployment, we employ a simple heuristic: after deployment, we merge with the contract having the largest bytecode size.  [^6] For proxy contracts, we use the sum of the bytecode sizes of all implementation contracts. For contracts implementing the ERC20 token standard (28), we discount the size contribution of ERC20 functions, following Heuristic 1 in Section 5.2. Larger contracts typically implement core protocol logic and are more likely to be frequently invoked by users.
Optimization 4: Bypassing Validation for Administrator Transactions. Transactions originated from administrators are exempted from control flow validation under the assumption that protocol administrators act benignly. In our evaluation in Section 6, we identify administrators as the deployers of protected contracts, allowing their transactions to bypass validation entirely. This optimization eliminates gas overhead for protocol maintenance operations while preserving security against external threats.